# Robust Control Lyapunov-Value Functions for Nonlinear Disturbed Systems

### v2

@@ A-A Proof of Proposition 2
Given any open set $\mathcal{C}$ and arbitrary control and disturbance signals, as long as the solution exists, the cost function
@@
For the local Lipschitzness in $x$, given any $x\text{ and }y\in\mathcal{C}$, $\epsilon>0$, from (10), we have
  1. 1.
    $\exists\lambda^{*}\in\Lambda,\forall u(\cdot)\in\mathbb{U}$,
    $V_{\gamma}(x,t)\leq J_{\gamma}(t,x,u(\cdot),\lambda^{*})+\epsilon,$
  2. 2.
    $\forall\lambda\in\Lambda$, $\exists u^{*}(\cdot)\in\mathbb{U}$,
    $V_{\gamma}(y,t)\geq J_{\gamma}(t,y,u^{*}(\cdot),\lambda)-\epsilon.$
@@
$V_{\gamma}(x,t)$ $-V_{\gamma}(y,t)$
$\leq$ $\bigl{(}J_{\gamma}(t,x,u^{*}(\cdot),\lambda^{*})+\epsilon\bigr{)}-\bigl{(}J_{\gamma}(t,y,u^{*}(\cdot),\lambda^{*})-\epsilon\bigr{)}$
$=$ $\bigl{(}J_{\gamma}(t,x,u^{*}(\cdot),\lambda^{*})-J_{\gamma}(t,y,u^{*}(\cdot),\lambda^{*})\bigr{)}+2\epsilon.$
Further, we have:
$\|J_{\gamma}(t,x,u^{*}(\cdot),\lambda^{*})-J_{\gamma}(t,y,u^{*}(\cdot),\lambda^{*})\|$
$=$ $\|\max_{s\in[t,0]}e^{\gamma(s-t)}\ell\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\big{)}-$
$\hskip 30.00005pt\max_{s\in[t,0]}e^{\gamma(s-t)}\ell\big{(}\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\big{)}\|$
$\leq$ $\max_{s\in[t,0]}\|e^{\gamma(s-t)}\ell\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\big{)}-$
$\hskip 30.00005pte^{\gamma(s-t)}\ell\big{(}\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\big{)}\|$
$\leq$ $\max_{s\in[t,0]}e^{\gamma(s-t)}\|\ell\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\big{)}-$
$\hskip 30.00005pt\ell\big{(}\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\big{)}\|$
$=$ $e^{-\gamma t}\|\ell\big{(}\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\big{)}-$
$\hskip 30.00005pt\ell\big{(}\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\big{)}\|.$
Plug in the definition of $\ell$, we have:
$\|J_{\gamma}(t,x,u^{*}(\cdot),\lambda^{*})-J_{\gamma}(t,y,u^{*}(\cdot),\lambda^{*})\|$
$\leq$ $e^{-\gamma t}\bigl{\|}\|\xi(s;t,x,u^{*}(\cdot),\lambda^{*})\|-\|\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\|\bigr{\|}$
$\leq$ $e^{-\gamma t}\|\xi(s;t,x,u^{*}(\cdot),\lambda^{*})-\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\|.$
$\|\xi(s;t,x,u^{*}(\cdot),\lambda^{*})-\xi(s;t,y,u^{*}(\cdot),\lambda^{*})\|\leq c\|x-y\|,$
Combined, we have
$\|V_{\gamma}(x,t)-V_{\gamma}(y,t)\|\leq e^{-\gamma t}c\|x-y\|+2\epsilon$ (39)
@@
$\hskip 60.00009pte^{\gamma(s-t_{1})}\ell(\xi(s;t_{1},x,u(\cdot),\lambda))$
@@
$V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max\biggl{\{}e^{\gamma(t_{1}-t)}V_{\gamma}(\xi(t_{1}),t_{1}),$
@@
$V_{\gamma}(x,t)\geq\sup_{\lambda\in\Lambda}\inf_{u\in\mathbb{U}}\max_{s\in[t,t_{1}]}e^{\gamma(s-t)}\ell(\xi(s)).$
@@
A-B Proof of Proposition 3
Since the convergence is point-wise, $\forall\epsilon>0$, $\exists t_{1},t_{2}<0$, s.t. $\forall x,y\in\mathcal{D}_{\gamma}$ we have
ϵitalic-$-\epsilon\leq$ $V_{\gamma}^{\infty}(x)-V_{\gamma}(x,t_{1})\leq\epsilon,$
ϵitalic-$-\epsilon\leq$ $V_{\gamma}^{\infty}(y)-V_{\gamma}(y,t_{2})\leq\epsilon.$
Since $V_{\gamma}(x,t)$ is non-increasing as $t\rightarrow-\infty$, taking $t_{N}=\min\{t_{1},t_{2}\}$, we have:
ϵitalic-$-\epsilon\leq$ $V_{\gamma}^{\infty}(x)-V_{\gamma}(x,t_{N})\leq\epsilon,$
ϵitalic-$-\epsilon\leq$ $V_{\gamma}^{\infty}(y)-V_{\gamma}(y,t_{N})\leq\epsilon.$
@@ A-B Proof of Proposition 3
$\|V_{\gamma}^{\infty}(x)-V_{\gamma}^{\infty}(y)\|\leq$ $\|V_{\gamma}(x,t_{N})-V_{\gamma}(y,t_{N})\|+2\epsilon$
$\leq$ $e^{-\gamma t_{N}}c\|x-y\|+2\epsilon.$
where we used Proposition 2 for the last inequality. Since $\epsilon$ can be chosen to be arbitrarily small, we conclude that the CLVF is locally Lipschitz in $\mathcal{D}_{\gamma}$ (refer to the proof of Theorem 3.2 of [13].) Further, given any compact subset of $\mathcal{D}_{\gamma}$, the $V_{\gamma}^{\infty}$ is Lipschitz in this compact subset.

### v3

@@ A-A Proof of Proposition 1
Given any open set $\mathcal{C}$ and arbitrary control and disturbance signals. Since the solution of (1) is guaranteed to exist given assumptions ($A_{1}$) and ($A_{2}$), the cost function
@@
For the local Lipschitzness in $x$, given any $x,y\in\mathcal{C}$, $u(\cdot)\in\mathbb{U},\lambda\in\Lambda$, we have:
$\|J_{\gamma}(t,x,u(\cdot),\lambda)-J_{\gamma}(t,y,u(\cdot),\lambda)\|$
$=$ $\|\max_{s\in[t,0]}e^{\gamma(s-t)}\ell\big(\xi(s;t,x,u(\cdot),\lambda)\big)$-
$\hskip 30.00005pt\max_{s\in[t,0]}e^{\gamma(s-t)}\ell\big(\xi(s;t,y,u(\cdot),\lambda)\big)\|$
$\leq$ $\max_{s\in[t,0]}\|e^{\gamma(s-t)}\ell\big(\xi(s;t,x,u(\cdot),\lambda)\big)$-
$\hskip 30.00005pte^{\gamma(s-t)}\ell\big(\xi(s;t,y,u(\cdot),\lambda)\big)\|$
$\leq$ $\max_{s\in[t,0]}e^{\gamma(s-t)}\|\ell\big(\xi(s;t,x,u(\cdot),\lambda)\big)$-
$\hskip 30.00005pt\ell\big(\xi(s;t,y,u(\cdot),\lambda)\big)\|$
$=$ $e^{-\gamma t}\|\ell\big(\xi(s;t,x,u(\cdot),\lambda)\big)$-
$\hskip 30.00005pt\ell\big(\xi(s;t,y,u(\cdot),\lambda)\big)\|.$
Plug in the definition of $\ell$, we have:
$\|J_{\gamma}(t,x,u(\cdot),\lambda)-J_{\gamma}(t,y,u(\cdot),\lambda)\|$
$\leq$ $e^{-\gamma t}\bigl\|\|\xi(s;t,x,u(\cdot),\lambda)\|-\|\xi(s;t,y,u(\cdot),\lambda)\|\bigr\|$
$\leq$ $e^{-\gamma t}\|\xi(s;t,x,u(\cdot),\lambda)-\xi(s;t,y,u(\cdot),\lambda)\|.$
Since this holds for arbitrary $u(\cdot)$ and $\lambda$, we have
$\|V_{\gamma}(x,t)-V_{\gamma}(y,t)\|$
$\leq$ $e^{-\gamma t}\|\xi(s;t,x,u(\cdot),\lambda)-\xi(s;t,y,u(\cdot),\lambda)\|.$
$\|\xi(s;t,x,u(\cdot),\lambda)-\xi(s;t,y,u(\cdot),\lambda)\|\leq c\|x-y\|,$
@@
$\|V_{\gamma}(x,t)-V_{\gamma}(y,t)\|\leq e^{-\gamma t}c\|x-y\|$ (48)
@@
$\hskip 60.00009pte^{\gamma(s-t_{1})}\ell(\xi(s;t_{1},x,u(\cdot),\lambda)).$
@@
$V_{\gamma}(x,t)=\sup_{\lambda\in\Lambda_{t}}\inf_{u(\cdot)\in\mathbb{U}_{t}}\max\biggl\{e^{\gamma(t_{1}-t)}V_{\gamma}(\xi(t_{1}),t_{1}),$
@@
$V_{\gamma}(x,t)\geq\sup_{\lambda\in\Lambda_{t}}\inf_{u(\cdot)\in\mathbb{U}_{t}}\max_{s\in[t,t_{1}]}e^{\gamma(s-t)}\ell(\xi(s)).$
@@
A-B Proof of Theorem 2
Proof:
Following [9],
  1. 1.
    $V_{\gamma}$ is a viscosity sub-solution of (21), if for any $\phi\in C^{1}(\mathbb{R}^{n}\times(-\infty,0])$, $(x_{0},t_{0})\in\mathbb{R}^{n}\times(-\infty,0]$ is a local maxima for $V_{\gamma}-\phi$, then
    $0\leq\max\biggl\{\ell(x_{0})-V_{\gamma}(x_{0},t_{0}),\hskip 10.00002ptD_{t}\phi(x_{0},t_{0})+$
    $\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\phi(x_{0},t_{0})\cdot f(x_{0},u,d)+\gamma V_{\gamma}(x_{0},t_{0})\biggl\},$ (49)
  2. 2.
    $V_{\gamma}$ is a viscosity super-solution of (30), if for any $\psi\in C^{1}(\mathbb{R}^{n}\times(-\infty,0])$, $(x,t)\in\mathbb{R}^{n}\times(-\infty,0]$ is a local minima for $V_{\gamma}-\psi$, then
    $0\geq\max\biggl\{\ell(x_{0})-V_{\gamma}(x_{0},t_{0}),\hskip 10.00002ptD_{t}\psi(x_{0},t_{0})+$
    $\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\psi(x_{0},t_{0})\cdot f(x_{0},u,d)+\gamma V_{\gamma}(x_{0},t_{0})\biggl\},$ (50)
W.L.O.G, we could always assume $x$ is a strict local maxima (minima), and the maximum (minimum) value is 0, i.e., $V_{\gamma}(x_{0},t_{0})-\phi(x,t_{0})=0$ ($V_{\gamma}(x_{0},t_{0})-\psi(x_{0},t_{0})=0$). Finally, $V_{\gamma}$ is a viscosity solution of (21) if it is both a sub-solution and a supersolution.
We will prove by contradiction. We start with the sub-solution. Assume (1) is wrong, then there exists $\theta_{1},\theta_{2}>0$ s.t. both the followings hold
$\ell(x_{0})-V_{\gamma}(x_{0},t_{0})\leq-\theta_{1}<0,$ (51)
$D_{t}\phi(x_{0},t_{0})+\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\phi(x_{0},t_{0})\cdot f(x_{0},u,d)$
$+\gamma V_{\gamma}(x_{0},t_{0})\leq-\theta_{2}<0.$ (52)
By continuity of $\ell$ and $\xi$, there exists $\delta_{1}>0$ s.t. for any $u(\cdot),\lambda[u]$, $s\in[t_{0},t_{0}+\delta_{1}]$
$|e^{\gamma(s-t_{0})}\ell(\xi(s;t_{0},x_{0},u,\lambda[u]))-\ell(x_{0})|\leq\frac{\theta_{1}}{2}.$ (53)
Combined with (51), we have
$e^{\gamma(s-t_{0})}\ell(\xi(s;t_{0},x_{0},u(\cdot),\lambda))$ $\leq\ell(x_{0})+\frac{\theta_{1}}{2},$
$\leq V_{\gamma}(x_{0},t_{0})-\frac{\theta_{1}}{2}.$ (54)
Further, since $V_{\gamma}(x_{0},t_{0})=\phi(x_{0},t_{0})$,  (52) can be written as
$D_{t}\phi(x_{0},t_{0})+\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\phi(x_{0},t_{0})\cdot f(x_{0},u,d)+$
$\gamma\phi(x_{0},t_{0})\leq-\theta_{2}$
and for any $d\in\mathcal{D}$,
$D_{t}\phi(x_{0},t_{0})+\min_{u\in\mathcal{U}}D_{x}\phi(x_{0},t_{0})\cdot f(x_{0},u,d)+\gamma\phi(x_{0},t_{0})$
$\leq-\theta_{2}.$
Since $\phi\in C^{1}$, there exists $\delta>0,\bar{u}\in\mathcal{U}$, $s\in[t_{0},t_{0}+\delta]$ s.t.
$D_{t}\phi(\xi(s),s)+D_{x}\phi(\xi(s),s)\cdot f(\xi(s),\bar{u}(s),\lambda[\bar{u}](s))$
$+\gamma\phi(\xi(s),s)\leq-\frac{\theta_{2}}{2}.$
Muliplly both side with $e^{\gamma(s-t_{0})}$ and integrate on both side for $s\in[t_{0},t_{0}+\delta]$, we get:
$e^{\gamma\delta}\phi(\xi(t_{0}+\delta;x_{0},t_{0},\bar{u}(\cdot),\lambda[\bar{u}]),t_{0}+\delta)-\phi(x_{0},t_{0})$
$\leq-\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$
Since $V_{\gamma}-\phi$ has local maxima at $(x_{0},t_{0})$ and the maximum value is 0, we have
$e^{\gamma\delta}V_{\gamma}(\xi(t_{0}+\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]),t_{0}+\delta)$
$\leq$ $\phi(x_{0},t_{0})-\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}=V_{\gamma}(x_{0},t_{0})-\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$ (55)
Combine (A-B) and (55), and because $d$ is arbitrary, we have
$V_{\gamma}(x_{0},t_{0})-\min\bigl\{\frac{\theta_{1}}{2},\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}\bigr\}$
$\geq$ $\sup_{\lambda}\max\bigl\{e^{\gamma(s-t_{0})}\ell(\xi(s;t_{0},x_{0},t,\lambda)),$
$\hskip 60.00009pte^{\gamma\delta}V_{\gamma}(\xi(t_{0}+\delta;x_{0},t_{0},\bar{u}(\cdot),\lambda[\bar{u}]),t_{0}+\delta)\bigr\}$
$\geq$ $V_{\gamma}(x_{0},t_{0}),$
where the last inequality is from (1). This is a contradiction, so (1) must hold, i.e., $V_{\gamma}$ is a viscosity sub-solution.
Next, we examine the super-solution. Assume (2) is wrong, then $\exists\theta_{1},\theta_{2}>0$ s.t. one of the followings hold
$\ell(x_{0})-V_{\gamma}(x_{0},t_{0})\geq\theta_{1}>0,$ (56)
$D_{t}\psi(x_{0},t_{0})+\max_{d\in\mathcal{D}}\min_{u\in\mathcal{U}}D_{x}\psi(x_{0},t_{0})\cdot f(x_{0},u,d)$
$+\gamma V_{\gamma}(x_{0},t_{0})\geq\theta_{2}>0.$ (57)
If (56) holds, then from (53), there exists $\delta_{1}>0$ s.t. for any $\theta_{1}>0,u(\cdot),\lambda$ and $s\in[t_{0},t_{0}+\delta_{1}]$
$e^{\gamma(s-t_{0})}\ell(\xi(s;t_{0},x_{0},u(\cdot),\lambda))$ $\geq\ell(x_{0})-\frac{\theta_{1}}{2}$
$\geq V_{\gamma}(x_{0},t_{0})+\frac{\theta_{1}}{2}.$
Plug in the DPP (1), we have
$V_{\gamma}(x_{0},t_{0})\geq$ $\sup_{\lambda}\inf_{u(\cdot)}\max_{s\in[t_{0},t_{0}+\delta]}e^{\gamma(s-t_{0})}\ell(\xi(s;t_{0},x_{0},u(\cdot),\lambda))$
$\geq$ $V_{\gamma}(x_{0},t_{0})+\frac{\theta_{1}}{2},$
which is a contradiction. Therefore (56) cannot be true. If (57) holds, then from the same derivation of (55), there exists $\delta_{2}>0$, $\exists\lambda,\forall u(\cdot)$, s.t. $\forall s\in[t,t+\delta_{2}]$
$e^{\gamma\delta}V_{\gamma}(\xi(t_{0}+\delta;t_{0},x_{0},\bar{u}(\cdot),\lambda[\bar{u}]),t_{0}+\delta)$
$\geq$ $\psi(x_{0},t_{0})+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}=V_{\gamma}(x_{0},t_{0})+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma}.$
Again plug in DPP (1), we have
$V_{\gamma}(x_{0},t_{0})\geq$ $\sup_{\lambda}\inf_{u(\cdot)}e^{\gamma\delta}V_{\gamma}(\xi(t_{0}+\delta;x,t,\bar{u}(\cdot),\lambda[\bar{u}]),t_{0}+\delta)$
$\geq$ $V_{\gamma}(x_{0},t_{0})+\frac{\theta_{2}(e^{\gamma\delta}-1)}{2\gamma},$
To prove the uniqueness, we apply the standard comparison principle. It should be noted that since we specify $\ell(x;p)=\|x-p\|-V^{\infty}_{m}$, the TV-R-CLVF (9) is radially unbounded. However, given any finite $t\leq 0$, its spatial growth is linearly bounded, i.e., there exists $C$ s.t. for all $x\in\mathbb{R}^{n},t\leq 0$
$V_{\gamma}(x,t)\leq C(1+\|x\|).$ (58)
This is because $f$ is bounded, and the exponential term is bounded by $e^{\gamma t}$. Consider continuous functions $V_{1}$ and $V_{2}$ both Lipschitz (with Lipschitz constant $L_{v}$), radially unbounded, and satisfying (58). Assume $V_{1}$ and $V_{2}$ are viscosity sub- and supersolutions of (21), respectively. Define
$V_{k}(x,t)=V_{1}(x,t)+k(a\|x\|^{2}+c)$
for constants $a,c,k>0$. Note that the growth of $V_{k}$ is quadratically bounded and Lipschitz, with Lipschitz constant $L_{k}$. Then $\ell(x)-V_{k}(x,t)<0$, and $V_{k}$ is the strict viscosity subsolution to
$D_{t}\phi+H(x,D_{x}\phi)+\gamma V_{k}\geq k,$
where $H$ is the Hamiltonian given by (22). Recall that $H$ is bounded above and Lipschitz continuous. Denote
$M_{k}=\sup_{[x,s]\in\mathbb{R}^{n}\times[t,0]}V_{2}(x,s)-V_{k}(x,s),$
and assume $M_{k}>0$. For any $\varepsilon,\eta,\delta>0$, define:
$\Phi_{\varepsilon,\delta}(x,y,s)=$ $V_{k}(x,s)-V_{2}(y,s)-\frac{\|x-y\|^{2}}{2\varepsilon}$
$-\delta(\|x\|^{4}+\|y\|^{4}).$ (59)
Since $\Phi\rightarrow$-$\infty$ as $\|x\|+\|y\|\rightarrow\infty$, $\Phi$ attains its maximum at some state $(\bar{x},\bar{y},\bar{s})$. Therefore, we have:
$\Phi(\bar{x},\bar{x},\bar{s})+\Phi(\bar{y},\bar{y},\bar{s})\leq 2\Phi(\bar{x},\bar{y},\bar{s}),$
which gives us
$\frac{\|\bar{x}-\bar{y}\|^{2}}{\epsilon}$ $\leq V_{k}(\bar{x},\bar{s})-V_{k}(\bar{y},\bar{s})+V_{2}(\bar{x},\bar{s})-V_{2}(\bar{y},\bar{s})$
$\leq(L_{k}+L_{v})\|\bar{x}-\bar{y}\|.$
@@ A-B Proof of Theorem 2
$\|\bar{x}-\bar{y}\|\leq(L_{k}+L_{v})\varepsilon,$ (60)
i.e., $\bar{x}-\bar{y}\rightarrow 0$ as $\varepsilon\rightarrow 0$. Define test functions
$\phi(x,t)$ $=V_{2}(\bar{y},\bar{s})+\frac{\|x-\bar{y}\|^{2}}{2\varepsilon}+\delta(\|x\|^{4}+\|\bar{y}\|^{4}),$
$\psi(y,t)$ $=V_{k}(\bar{x},\bar{s})-\frac{\|\bar{x}-y\|^{2}}{2\varepsilon}-\delta(\|\bar{x}\|^{4}+\|y\|^{4}),$
and we can check $V_{k}-\phi$ attains maximum at $(\bar{x},\bar{y},\bar{s})$, $V_{2}-\psi$ attains minimum at $(\bar{x},\bar{y},\bar{s})$. Since $V_{k}$ is a strict viscosity subsolution, the following hold:
$D_{t}\phi(\bar{x},\bar{s})+H(\bar{x},D_{x}\phi(\bar{x},\bar{s}))+\gamma V_{k}(\bar{x},\bar{s})\geq k.$ (61)
Since $V_{2}$ is a viscosity supersolution, from (2), both of the following hold
$\ell(\bar{y})-V_{2}(\bar{y},\bar{s})\leq 0,$ (62)
$D_{t}\psi(\bar{y},\bar{s})+H(\bar{y},D_{x}\psi(\bar{y},\bar{s}))+\gamma V_{2}(\bar{y},\bar{s})\leq 0.$ (63)
Then since $D_{t}\phi(\bar{x},\bar{s})=D_{t}\psi(\bar{y},\bar{s})$, from (61) and (63), we have
$\gamma(V_{2}(\bar{y},\bar{s})-V_{k}(\bar{x},\bar{s}))$
$\leq$ $-k+(H(\bar{x},D_{x}\phi(\bar{x},\bar{s}))-H(\bar{y},D_{x}\psi(\bar{y},\bar{s}))),$
$\leq$ $-k+L_{Hx}(\bar{x}-\bar{y})+L_{Hp}(D_{x}\phi(\bar{x},\bar{s})-D_{x}\psi(\bar{y},\bar{s})),$
where $L_{Hx}$ and $L_{Hp}$ are the Lipschitz constants of $H$ over $x$ and $p$ respectively. By the definition of $\phi,\psi$, we have
$D_{x}\phi(\bar{x},\bar{s})-D_{x}\psi(\bar{y},\bar{s})=3\delta(\|\bar{x}]\bar{x}-\|\bar{y}\|\bar{y}).$
Combining with (60), as $k\rightarrow 0$, we can always pick $\varepsilon,\delta$ small enough s.t.
$V_{2}(\bar{y},\bar{s})-V_{k}(\bar{x},\bar{s})\leq-\frac{k}{2}.$
This contradicts the assumption that $M_{k}>0$. Therefore, we proved $V_{2}\leq V_{1}$. Switching the role of $V_{1}$ and $V_{2}$ and following the identical process, we can prove $V_{1}\leq V_{2}$. Therefore the TV-R-CLVF has a unique viscosity solution.